Thursday, April 14, 2011

Bad-weather testing

After some post-lift fixing & de-bugging, Tech Ops reckons they've got the RSS filter magazine & slit mask problems sorted out - so we're good to go!  They left us with the following warning last night...

Sadly though, the weather's been Anything but co-operative.  A dramatic (i.e. astronomically unpretty) sunset was followed by more rain & a feeble lightning display - so Ken was restricted to closed-dome tests such as focus runs & stray light observations.

Today the guys got the etalons connected up & tested (that all looks fine), various blocked/dirty filters & valves were sorted out to eliminate overheating problems & they also worked on some RSS baffling between the filters & the detector.  However, the undisputed highlight of the day was a display by *3* adult black eagles over the plateau, just a pity they were too far away to photograph properly...

Alas - once more, a beautiful clear morning brewed up into a cloudy afternoon & evening with rain + heavy cloud about, so we're likely to be playing the waiting game again tonight :(

No comments:

Post a Comment